# TaCo: Targeted Concept Removal in Output Embeddings for NLP via Information Theory and Explainability

### v1

@@ Abstract
The fairness of Natural Language Processing (NLP) models has emerged as a crucial concern. Information theory indicates that to achieve fairness, a model should not be able to predict sensitive variables, such as gender, ethnicity, and age. However, information related to these variables often appears implicitly in language, posing a challenge in identifying and mitigating biases effectively.
To tackle this issue, we present a novel approach that operates at the embedding level of an NLP model, independent of the specific architecture. Our method leverages insights from recent advances in XAI techniques and employs an embedding transformation to eliminate implicit information from a selected variable. By directly manipulating the embeddings in the final layer, our approach enables a seamless integration into existing models without requiring significant modifications or retraining.
In evaluation, we show that the proposed post-hoc approach significantly reduces gender-related associations in NLP models while preserving the overall performance and functionality of the models.
An implementation of our method is available: https://github.com/fanny-jourdan/TaCo
@@ 1 Introduction
The multiple and impressive successes of Artificial Intelligence and Machine Learning have made them more and more ubiquitous in applications where a model’s decisions has genuine repercussions on people’s lives. These decisions, however, typically come with biases that can lead to unfair, unethical consequences. To guard against such consequences, we must identify and understand the biased information these models utilize. This links the task of ensuring an ethical comportment of a model with its interpretability or explainability: if we can show the real reasons or features why a model makes the predictions or recommendations it does, then we can control for biases among these features.
In this paper, we propose a method for finding factors that influence model predictions using an algebraic decomposition of the latent representations of the model into orthogonal dimensions. Using the method of Singular Value Decomposition (SVD), which is well suited to our particular test case, we isolate various dimensions that contribute to a prediction. As our case study, we study the use of gender information in predicting occupations in the Bios dataset De-Arteaga et al. (2019) by a RoBERTa model Liu et al. (2019a). By intervening on the latent representations to remove certain dimensions and running the model on the edited representations, we show first a causal connection between various dimensions and model predictions about gender and separately about occupation. This enables us then to eliminate dimensions in the decomposition that convey gender bias in the decision without much affecting the model’s predictive performance on occupations. Because SVD gives us orthogonal dimensions, we know that once we have removed a dimension, the information it contains is no longer in the model.
Crucially, we show how to link various dimensions in the SVD decomposition of the latent representation with concepts like gender and its linguistic manifestations that matter to fairness, providing not only a faithful explanation of the model’s behaviour but a humanly understandable or plausible one as well. Faithfulness provides a causal connection and produces factors necessary and sufficient for the model to produce its prediction given a certain input. Plausibility has to do with how acceptable the purported explanation is to humans Jacovi and Goldberg (2020).
In addition, we demonstrate that our method can reach close to the theoretically ideal point of minimizing algorithmic bias while preserving strong predictive accuracy, and we offer a way of quantifying various operations on representations. Our method also has a relatively low computational cost, as we detail below. In principle, our approach applies to any type of deep learning model; it also can apply to representations that are generated at several levels, as in modern transformer architectures with multiple attention layers Vaswani et al. (2017). As such it can serve as a diagnostic tool for exploring complex deep learning models.
Figure 1: Overview of our method. The SVD decomposition (I) yields concepts, whose importance (II) with respect to gender and occupation are evaluated with Sobol method. Finally, some concepts are removed, which “neutralizes” the gender information and (III) produces a fairer classifier.
@@ 2 Related work
In this section, we review pertinent literature to the two strands of our paper, bias and explainability.
2.1 Gender Bias in Natural Language Processing models
In this paper, we aim to enhance the fairness of our model by employing a prevalent approach in information theory, which involves eliminating gender-related information to prevent the model from utilizing it Kilbertus et al. (2017). This method of achieving gender fairness has its own body of literature:
@@ 2.1 Gender Bias in Natural Language Processing models
A simple approach in the literature is to remove explicit gender indicators (he, she, her, etc.) directly from the dataset De-Arteaga et al. (2019). However, modern transformer models have been shown to be able to learn implicit gender indicators Devlin et al. (2018); Karita et al. (2019). This essentially limits the impact that such a naïve approach can have on this kind of models. As a matter of fact, a RoBERTa model trained without these gender indicators can still predict with upwards of 90% accuracy the gender of a person writing a short LinkedIn bio (see Appendix A).
@@
Another approach attempts to create directly word embeddings that do not exhibit gender bias Bolukbasi et al. (2016); Garg et al. (2018); Zhao et al. (2018a, b); Caliskan et al. (2017). Although interesting in principle, Gonen and Goldberg (2019) shows these methods may camouflage gender biases in word embeddings but fail to eliminate them entirely.
In addition, gender information is usually correlated with other features and manifests itself at different levels in the sentence’s structure, effectively leaking through in the shape of implicit bias that can be easily exploited by transformer models to generate biased predictions.
A third approach, and more in line with our proposed method, consists in seeking fairness in the sense of information theory Xu et al. (2020). Specifically, we propose to use methodology from the field of explainable AI (XAI) to perform feature selection to reduce the effect of algorithmic bias Frye et al. (2020); Dorleon et al. (2022); Galhotra et al. (2022) in a human-understandable yet principled way. Indeed, our aim is to decompose the model’s embedding into an interpretable base of concepts – i.e. our features –, select those that carry the most information about gender and the least about the actual task, and suppress them from the embedding.
2.2 Explainability and Interpretability in Model Analysis
Explainability
We therefore employ explainability methods to achieve fairness in this context. Let us first provide an overview of this field.
Classical approaches analyse how each part of the input influences the model’s prediction, either through perturbation Ribeiro et al. (2016); Zeiler and Fergus (2014) or by leveraging the gradients inside the neural network Sundararajan et al. (2017). However, these methods are prone to adversarial manipulation Wang et al. (2020), perform only partial input recovery Adebayo et al. (2018), and exhibit a general lack of stability with respect to the input Ghorbani et al. (2019). Moreover, they often fail to provide full interpretability, because they provide explanations that tamper with some features but not others; in general the features tampered with do not offer a guarantee of the outcome Ignatiev et al. (2019); Marques-Silva (2023).
Rationalization approaches also play a pivotal role in enhancing our understanding Bouchacourt and Denoyer (2019); Antognini and Faltings (2021).
Asher et al. (2021) explored counterfactual approaches to isolating features that may have harmful biases. Another branch of explicability close to this one is contrastive explanations, recently used in Yin and Neubig (2022); Ferrando et al. (2023) .
The most recent approaches are concept-based methods such as TCAV Kim et al. (2018) and CRAFT Fel et al. (2023) for vision, and COCKATIEL Jourdan et al. (2023) for NLP.
Faithfulness and plausibility
When talking of explanations, we will speak of the explanans, the explaining part, and the explanandum, the part to be explained. Discussions of explainability have led to different interpretations of explainability and interpretability. In this section, we set out how we understand these terms. In the recent literature they have sometimes been used equivalently Jacovi and Goldberg (2020), but also sometimes distinguished Rudin (2019) using the concepts of faithfulness and plausibility. Plausibility has to do with how acceptable the purported explanation is to humans Jacovi and Goldberg (2020). Faithfulness, on the other hand, involves a causal connection (what is causally necessary and sufficient for the model to produce its prediction given a certain input); a faithful explanatory model is one that captures these causal relations in the explanations it generates.
To capture these causal relations, we will perform interventions removing the putative causal factors to see how this affects the model’s output. What intuitively needs to be done is an intervention that removes the impact of a certain token from the model’s prediction pipeline. However, simply masking a token from another is not sufficient to do this, because all tokens interact at multiple levels in modern transformer architectures. Blocking their interaction at one level does not block interactions at later or higher levels that can affect downstream predictions . Thus, trying to construct faithful models and causal connections requires some care.
@@ 3 Preliminaries and definitions
3.1 Methodology
The idea behind the method we propose for creating gender-neutral embedding can be summed up in 3 components: (i) Identification of occupation-related concepts within the latent space of an NLP model; (ii) evaluation of the significance of each concept in predicting gender; and (iii) removal of specific concepts to construct a gender-neutral embedding, followed by retraining a classifier to predict occupation based on this modified embedding. For a visual representation of the method, refer to Figure 1, which provides an explanatory diagram illustrating the different stages of the process.
To realize the proposed method, it is crucial to secure certain guarantees and engage in rigorous theoretical modelling. This ensures the identification and application of suitable, practical tools. This involves establishing clear criteria for selecting practical tools and methodologies, enabling the achievement of dependable and consistent results. These are defined in this section.
@@ 3.2 Notation
With $n$ individuals in the training set, we define $\bm{X}=(\bm{x}_{1},...,\bm{x}_{n})\in\mathcal{X}^{n}$, $\bm{Y}=(y_{1},...,y_{n})\in\mathcal{Y}^{n}$, and $\bm{G}=(g_{1},...,g_{n})\in\{\text{Male, Female}\}^{n}$.
The model $\bm{f}$ is conceptualized as a composition of two functions. First, $h$ encompasses all layers from the input $\bm{x}$ to a latent space, which is a transformed representation of the input data. Then, the classification layers $c$ classify these transformed data. The model can be summarized as:
$\bm{f}(\bm{x})=c\circ h(\bm{x})\,,$ (1)
where $h(\bm{x})\in\mathbb{R}^{d}$. In this space, we obtain an output embedding matrix, denoted as $\bm{A}=h(\bm{X})\in\mathbb{R}^{n\times d}$, which is also referred to as the CLS embedding in the context of an encoder transformer for classification task.
3.3 Causal underpinnings of the method
A faithful reconstruction of $\bm{f}$ should support a causal relation that can be formulated as follows: $A$ and $B$ and had $A$ not been the case, $B$ would not have been the case either Lewis (1973); Pearl (2009); Jacovi and Goldberg (2020). $A$ describes the causally necessary and “other things being equal” sufficient condition for $B$. Faithfulness implies that a particular type of input is causally necessary for the prediction at a given data point $\bm{x}$. Such a condition could take the form of one or more variables having certain values or set of variables. It can also take the form of a logical statement at a higher degree of abstraction.
A counterfactual theory is simply a collection of such statements, and for every deep learning model, there is a counterfactual theory that describes it Jacovi et al. (2021); Asher et al. (2022); Yin and Neubig (2022).
[…]
For instance, we can alternatively choose (i) $\mathscr{W}=\mathcal{X}^{n}$ the model’s input entries, (ii) $\mathscr{W}=$ the set of possible attention weights in $\bm{f}$ (assuming it’s a transformer model), (iii) $\mathscr{W}=$ the possible parameter settings for a specific intermediate layer of $\bm{f}$ –e.g., the attribution matrix as in Fel et al. (2023), (iv) $\mathscr{W}=$ the possible parameter settings for the final layer of an attention model as suggested in Wiegreffe and Pinter (2019), or (iv) sets of factors or dimensions of the final layer output matrix from the entire model Jourdan et al. (2023). Our explanatory model here follows the ideas of Jourdan et al. (2023), though it differs importantly in details. For one thing, we note that our general semantic approach allows us to have Boolean combinations even first order combinations of factors to isolate a causally necessary and sufficient explanans.
@@ 3.3.1 Counterfactual intervention for concept ranking
In part 2 of our method, we calculate the causal effect on the sensitive variable for each dimension $r$ of $\bm{U}$ and $\bm{W}$ of the $\bm{A}=\bm{U}\bm{W}$ decomposition created in part 1 of the method (more details on this decomposition in practice in section 4.1). We conduct a counterfactual intervention that involves removing the dimension or “concept” $r$. By comparing the behaviour of the model with the concept removed and the behaviour of the original model, we can assess the causal effect of the concept on the sensitive variable. As explained above, this assumes a counterfactual model where $\mathscr{W}$ corresponds to the possible variations of $\bm{A}=h(\bm{X})$ when reducing the dimensionality of its decomposition $\bm{A}=\bm{U}\bm{W}$.
In addition, this approach requires that we are really able to remove the targeted dimension, which means that the dimensions must be independent of each other.
@@ 3.4 Information theory guarantees
After performing our counterfactual intervention, we need to compute the "importance of a dimension" for the sensitive variable. Our goal is for our final embedding $\bm{A}$ to be as independent as possible from the sensitive variable $\bm{G}$.
According to information theory, if we are unable to predict a variable $\bm{G}$ from the data $\bm{A}$, it implies that this variable $\bm{G}$ does not influence the prediction of $y$ by $\bm{A}$ Xu et al. (2020). This definition is widely used in causal fairness studies Kilbertus et al. (2017), making it a logical choice for our task.
We train a classifier $c_{g}$ for the gender $\bm{G}$ prediction from the matrix $\bm{A}$ (see appendix C.2 for details on its architecture). If $c_{g}$ cannot achieve high accuracy for gender prediction, it implies that $\bm{A}$ does not contain gender-related information. The most reliable approach is thus to remove each of the dimensions found by the decomposition one by one and re-train a classifier to predict gender from this matrix. This can be computationally expensive if we have many dimensions. An alternative is to train only one classifier to predict gender on the original matrix $\bm{A}$, and then use a sensitivity analysis method on the dimensions to determine their importance for prediction. We will choose the latter for our method. However, we will still demonstrate the relevance of this approach by training a classifier each time we remove a dimension and show its effect in Section 5.2.

### v4

# TaCo: Targeted Concept Erasure Prevents Non-Linear Classifiers From Detecting Protected Attributes

@@ Abstract
Ensuring fairness in NLP models is crucial, as they often encode sensitive attributes like gender and ethnicity, leading to biased outcomes. Current concept erasure methods attempt to mitigate this by modifying final latent representations to remove sensitive information without retraining the entire model. However, these methods typically rely on linear classifiers, which leave models vulnerable to non-linear adversaries capable of recovering sensitive information.
We introduce Targeted Concept Erasure (TaCo), a novel approach that removes sensitive information from final latent representations, ensuring fairness even against non-linear classifiers. Our experiments show that TaCo outperforms state-of-the-art methods, achieving greater reductions in the prediction accuracy of sensitive attributes by non-linear classifier while preserving overall task performance. Code is available on GitHub.
@@ 1 Introduction
Artificial Intelligence models are increasingly deployed in critical domains but often unintentionally encode sensitive attributes such as gender, ethnicity, and age, leading to unfair and unethical outcomes. Ensuring fairness in these models is imperative; and according to information theory, a fair model should be incapable of predicting these sensitive variables, regardless of whether the predictor is linear or non-linear.
Concept erasure methods have been developed to meet this desideratum. These methods modify the representation in the latent space of NLP models just before classification, attempting to remove information about the sensitive variable in this representation, to enable fairer classification without the need to re-train the entire model. These methods aim to hide sensitive information in final latent space of models, but generally evaluate their effectiveness using linear classifiers such as logistic regression (Ravfogel et al., 2020, 2022; Belrose et al., 2023). While these approaches reduce the ability of linear models to recover sensitive information, they often fail to eliminate it entirely, leaving it accessible to non-linear models (Ravfogel et al., 2023). This is a significant limitation for many concept erasure methods to achieve true fairness.
To address this limitation, we propose Targeted Concept Erasure (TaCo), a novel concept erasure method that effectively removes sensitive information from final latent representations, ensuring fairness even against non-linear adversaries. Our approach involves three key steps:
(i) Concept Discovery: We identify relevant concepts within the latent space that are associated with the sensitive attribute.
(ii) Concept Ranking: We rank these concepts based on their importance to both the sensitive variable and the target label. This ranking allows us to find the optimal trade-off between fairness and task performance.
(iii) Concept Erasure: We systematically remove concepts, prioritizing those that are most influential to the sensitive attribute but have minimal impact on the primary task prediction.
We evaluate our method on the Bios dataset (De-Arteaga et al., 2019) using a variety of pre-trained language models, including RoBERTa (Liu et al., 2019a), DistilBERT (Sanh et al., 2019), DeBERTa (He et al., 2020), and T5 (Raffel et al., 2020). Our experiments demonstrate that TaCo significantly outperforms state-of-the-art concept erasure methods in preventing non-linear classifiers, such as multilayer perceptrons (MLPs), from predicting sensitive attributes. Specifically, TaCo achieves a greater reduction in the accuracy of predicting sensitive variables across all tested models, without compromising overall (intended) task performance.
In addition to enhancing fairness, our method offers interpretability benefits. By identifying and manipulating specific concepts, we answer the question: "Which concepts have an influence?". And we further extend our approach by applying concept-based explainability techniques –specifically, COCKATIEL (Jourdan et al., 2023)– directly on the identified concepts. This allows us to also answer the question: "Why did this concept have an influence?", providing deeper understanding and transparency into the model’s decision-making process.
@@ 2 Related work
In information theory, a common approach to improving the fairness of a model is to eliminate information about sensitive variables to prevent the model from using it (Kilbertus et al., 2017). This approach to fairness, which we’ll call "sensitive variable information erasure", has already been used in various forms in the literature
@@
One approach just removes explicit gender indicators (he, she, her, etc.) directly from the dataset (De-Arteaga et al., 2019). However, modern transformer models are able to learn implicit gender indicators (Devlin et al., 2018; Karita et al., 2019), which essentially limits the impact that such a naive approach can have on state-of-the-art models. In fact, a RoBERTa model trained without these gender indicators can still predict with upwards of 90% accuracy the gender of a person writing a short LinkedIn bio (see Appendix A).
@@
A second approach attempts to create word embeddings that do not exhibit gender bias (Bolukbasi et al., 2016; Garg et al., 2018; Zhao et al., 2018a, b; Caliskan et al., 2017; Bolukbasi et al., 2016; Zhao et al., 2019). Although interesting in principle, Gonen and Goldberg (2019) shows these methods may camouflage gender biases in word embeddings, but fail to eliminate them entirely.
A third approach, more in line with our proposed method, consists in intervening on the model’s latent space. Liang et al. (2020) introduce a technique based on decomposing the latent space using a PCA and subtracting from the inputs the projection into a bias subspace that carries the most information about the target concept. Another vein focuses on modifying the latent space with considerations coming from the field of information theory (Xu et al., 2020). Indeed, the $\nu$-information – a generalization of the mutual information – has been used as a means to carefully remove from the model’s latent space only the information related to the target concept. In particular, popular approaches include the removal of information through linear adversarials on the last layer of the model (Ravfogel et al., 2020, 2022), through Assignment-Maximization adversarials (Shao et al., 2023), and through linear models that act on the whole model (Belrose et al., 2023). However, Ravfogel et al. (2023) has recently shown a limitation of these methods. The limitation comes from the practical use of the $\nu$-information ($\nu$-information is a metric that updates the mutual information, considering the expressivity of functions measuring information in the latent space). Ravfogel et al. (2023) explains that by protecting information using a linear model, more expressive models can still decode this protected information. Our method avoids this problem, because we are effectively destroying information – and not simply hiding it from linear models. Thus, no model can access it.
In addition, our method, inspired by concept-based explainability techniques like TCAV (Kim et al., 2018), CRAFT (Fel et al., 2023), and COCKATIEL (Jourdan et al., 2023), provides the added benefit of enabling interpretability of the erased concepts. Unlike most concept erasure methods, which focus solely on removing sensitive information, our approach offers insight into how these concepts influence model predictions, enhancing transparency and interpretability without sacrificing performance.
Figure 1: Overview of TaCo method. A decomposition of the final latent embedding matrix (I) yields concepts, whose importance (II) with respect to the sensitive variable and the label are evaluated with Sobol method. Finally, some concepts are removed, which “neutralizes” the sensitive variable information and (III) produces a fairer classifier.
@@ 3 Preliminaries and definitions
The idea behind our method for concept erasure has 3 components: (i) identifies label-related concepts within the final latent space of an NLP model (such as the [CLS] embedding for encoder transformers like BERT, the EOS token embedding for encoder-decoder transformers like T5, or the first predicted token for decoder transformers like Llama). (ii) evaluates the importance of each concept for the prediction of the sensitive variable and the label to find the concepts to remove with the best trade-off between removing sensitive information and maintaining model performance; (iii) removes specific concepts to construct a new final latent embedding without sensitive variable information, after which we retrain a classifier to predict label based on this modified embedding. For a visual representation of the method, refer to Figure 1, which provides an explanatory diagram illustrating the different stages of the process.
Our method relies on a mathematical foundation that guarantees removal and ensures the identification and application of suitable, practical tools. This foundation also establishes clear criteria for selecting practical tools and methodologies, enabling the achievement of dependable and consistent results. These are defined in the next section.
@@ 3.1 Notation
We work in context of supervised learning assume that a neural network model, denoted as $\bm{f}\colon\mathcal{X}\rightarrow\mathcal{Y}$, has been previously trained to perform a specific classification task. Here, $\bm{x}\in\mathcal{X}$ represents an embedded input text (in our example, a LinkedIn biography), and $y\in\mathcal{Y}$ denotes their corresponding label (in our example, the occupation). We introduce an additional set, $g\in\mathcal{G}$, representing the sensitive variables (in our example, the gender). This variable is crucial as it is the one we have to control to ensure fairness in our model. With $n$ individuals in the training set, we define $\bm{X}=(\bm{x}_{1},...,\bm{x}_{n})\in\mathcal{X}^{n}$, $\bm{Y}=(y_{1},...,y_{n})\in\mathcal{Y}^{n}$, and $\bm{G}=(g_{1},...,g_{n})\in\mathcal{G}^{n}$.
We represent the model $\bm{f}$ as a composition of two functions.
$\bm{f}(\bm{x})=c\circ h(\bm{x})\,,$ (1)
where $h(\bm{x})\in\mathbb{R}^{d}$. $h$ encompasses all layers from the input $\bm{x}$ to a latent space, which is a transformed representation of the input data. The classification layers $c$ classify these transformed data. $h(\bm{X})$ is the final latent embedding matrix, $\bm{A}=h(\bm{X})\in\mathbb{R}^{n\times d}$, which is also referred to as the [CLS] embedding in the context of an encoder transformer, as the EOS token embedding for encoder-decoder transformers and the first predicted token for decoder transformers for a classification task.
3.2 Causal underpinnings of the method
A faithful reconstruction of $\bm{f}$ should support a causal relation: $A$ and $B$ and had $A$ not been the case, $B$ would not have been the case either (Lewis, 1973; Pearl, 2009; Jacovi and Goldberg, 2020). $A$ describes the causally necessary and “other things being equal” sufficient condition for $B$. Faithfulness implies that a particular input type is causally necessary for the prediction at a given data point $\bm{x}$. Such a condition could involve specific variable values or a set of variables and can also be a higher-level logical statement.
@@ 3.2 Causal underpinnings of the method
For example, we can choose: (i) $\mathscr{W}=\mathcal{X}^{n}$ (model’s input entries), (ii) $\mathscr{W}$ as the set of possible attention weights in $\bm{f}$ (if it’s a transformer model), (iii) $\mathscr{W}$ as the possible parameter settings for a specific intermediate layer of $\bm{f}$ (e.g., attribution matrix as in (Fel et al., 2023)), (iv) $\mathscr{W}$ as the possible parameter settings for the final layer of an attention model as suggested in (Wiegreffe and Pinter, 2019), or (v) sets of factors or dimensions of the final layer embedding matrix from the entire model (Jourdan et al., 2023).
Our explanatory model follows the ideas of (Jourdan et al., 2023) but differs in details. Notably, our semantic approach allows Boolean combinations, even first-order combinations of factors, to isolate a causally necessary and sufficient explanans.
@@ 3.2.1 Counterfactual intervention for concept ranking
In part 2 of our method, we calculate the causal effect on the sensitive variable for each dimension $r$ of $\bm{U}$ and $\bm{W}$ of the $\bm{A}=\bm{U}\bm{W}$ decomposition created in part 1 of the method (for more details on this decomposition see section 4.1). We conduct a counterfactual intervention that involves removing the dimension or “concept” $k\in[\![1;r-1]\!]$. By comparing the behavior of the model with the concept removed and the behavior of the original model, we can assess the causal effect of the concept on the sensitive variable. As explained above, this assumes a counterfactual model where $\mathscr{W}$ corresponds to the possible variations of $\bm{A}=h(\bm{X})$ when reducing the dimensionality of its decomposition $\bm{A}=\bm{U}\bm{W}$.
To really able to remove the targeted dimension, the dimensions must be orthogonal of each other. Our mathematical framework guarantees this as far as possible.
@@ 3.3 Information theory guarantees
As part of our counterfactual intervention, we calculate the "importance of a dimension" for the sensitive variable, so that our final integration $\bm{A}$ is as independent as possible from the sensitive variable $\bm{G}$. According to information theory, if we are unable to predict a variable $\bm{G}$ from the data $\bm{A}$, it implies that this variable $\bm{G}$ does not influence the prediction of $y$ by $\bm{A}$ (Xu et al., 2020). This definition is widely used in causal fairness studies (Kilbertus et al., 2017), making it a logical choice for our task.
To evaluate the independence between $\mathbf{a}$ and $\mathbf{g}$, we train a classifier $c_{g}$ to predict the sensitive variable $\mathbf{g}$ from the representation $\mathbf{a}$ (see Appendix C.2 for architectural details). If $c_{g}$ cannot achieve high accuracy in predicting $\mathbf{g}$, it suggests that $\mathbf{a}$ lacks information about the sensitive attribute.
While existing methods often employ linear classifiers to estimate this predictability, focusing solely on linear dependencies may overlook non-linear relationships, allowing residual sensitive information to remain accessible to non-linear models (Ravfogel et al., 2023). To address this limitation, we utilize a non-linear classifier—specifically, a multilayer perceptron (MLP)—to estimate the predictability of $\mathbf{g}$ from $\mathbf{a}$. This approach enables us to detect and mitigate both linear and non-linear dependencies, providing a more robust assessment of the residual sensitive information.
Calculating the importance of each dimension by individually removing them and retraining the classifier can be computationally intensive, especially in high-dimensional spaces. To circumvent this, we train a single non-linear classifier on the original representation $\mathbf{a}$ and apply a sensitivity analysis method (Section 4.2) to determine each dimension’s importance for predicting $\mathbf{g}$. This strategy efficiently identifies the dimensions most influential to the sensitive attribute without the need for exhaustive retraining.
For quantifying the degree of dependence between $\mathbf{a}$ and $\mathbf{g}$, we adopt the $\nu$-information metric. While this metric has been well-established in prior work using linear classifiers (Ravfogel et al., 2023; Belrose et al., 2023), we extend its application by incorporating non-linear classifiers in its computation. This extension allows us to capture non-linear dependencies that linear models might miss, offering a more comprehensive measure of mutual information.
By using a non-linear classifier to compute $\nu$-information, we provide a stronger guarantee of fairness against adversaries employing complex, non-linear models. In Section 5.2, we show that our method significantly reduces the presence of sensitive information, as evidenced by the decreased predictability of $\mathbf{g}$ even when using sophisticated non-linear classifiers.